# Video-Language Critic: Transferable Reward Functions for Language-Conditioned Robotics

### v1

@@ 6 Conclusion
We proposed Video-Language Critic (VLC), an approach to pretrain a foundation reward model for vision-language manipulation. In particular, we train our critic using contrastive video-language alignment and a ranking loss encouraging monotonic increases for successful trajectories. Our model predicts a language-conditioned state-value function conditioned on only a history of image observations, and can therefore readily be scaled to leverage external observation data from other actors. VLC can be used for various downstream tasks, such as model-free and model-based reinforcement learning. Further, we experimentally validated its usefulness as a reward function in out-of-distribution tasks (unseen, held-out environments in the same domain) and out-of-domain tasks (an unseen environment and embodiment in a different domain). Our experiments on Meta-World demonstrated improved results compared to 5 prior methods and a sparse reward baseline, with success rate increases of 12 and 34 percentage units, respectively, and sample efficiency gains of 65% and 300%, respectively. Unlike methods based on predicting expert actions in a given embodiment, VLC remains applicable in the absence of action labels. Therefore, it can be extended to equally learn from videos of humans in future work.
Acknowledgments and Disclosure of Funding
@@ B.2 Policy training
Training length and hence computational cost varies considerably across tasks. We terminate training after convergence to $\geq$98success (averaged over the 10 most recent evaluations), or after a maximal training length set per task. The resulting average training length across all 25 tasks was 600k environment steps for VLC MW50 and 750k for the Open X trained model. The corresponding GPU hours vary slightly based on exact architecture used, but we obtained approximately 750k steps in 24 hours on a single NVIDIA V100 or P100 GPU. Our total computational budget was therefore 19.2 GPU hours $\times$ 5 random seeds $\times$ 25 tasks = 2,400 GPU hours for the RL training experiments with VLC MW50 and 24 $\times$ 5 $\times$ 25 = 3,000 hours with VLC Open X. For the MW40 task generalization experiments, the average training length was again 600k for VLC, so the corresponding cost for training on 6 tasks was approximately 580 GPU hours for our method, and slightly more for baselines that took longer to train.
Appendix C Baseline implementations
For the LOReL baseline experiment, we used the proposed binary classification objective, reversed negatives and 2-frame conditioning (first and last), while keeping all of our other training details and data augmentations identical to our method. This is to ensure LOReL’s smaller and older original architecture did not account for any difference in performance. The original implementations were used for all other methods, and pretrained checkpoints were reused when applicable (RoboCLIP was pretrained on HowTo100M [24], LIV on EPIC-KITCHENS [8] and Voltron on Something-Something [11]).
@@ Appendix C Baseline implementations
Appendix D Meta-World environment definitions

### v2

@@ 5 Conclusion
We proposed Video-Language Critic (VLC), a method for training a foundation reward model for vision-language manipulation. In particular, we train our critic using contrastive video-language alignment and a ranking loss encouraging monotonic increases for successful trajectories. Our model predicts a language-conditioned state-value function conditioned on only a history of image observations, and can therefore readily be scaled to leverage external observation data from other actors. VLC can be used for various downstream tasks, such as model-free and model-based reinforcement learning. Further, we experimentally validated its usefulness as a reward function in out-of-distribution tasks (unseen, held-out environments in the same domain) and out-of-domain tasks (an unseen environment and embodiment in a different domain). Our experiments on Meta-World demonstrated improved results compared to 5 prior methods and a sparse reward baseline, with success rate increases of 12 and 34 percentage units, respectively, and sample efficiency gains of 65% and 300%, respectively. Unlike methods based on predicting expert actions in a given embodiment, VLC remains applicable in the absence of action labels. Therefore, it can be extended to equally learn from videos of humans in future work.
Acknowledgments
@@ B.2 Policy training
Reward normalization
As reward model predictions for the starting state can vary across initial states, even for the same task, we shift the rewards of the episode such that $r_{1}=0$, as the subsequent behavior should be scored relative to this state. In addition to this, we apply the standard normalization logic from Gymnasium [46], which scales reward values such that their exponential moving average has fixed variance $(1-\gamma)^{2}$.
Reward component weighting
@@
Computational cost
Training length and hence computational cost varies considerably across tasks. We terminate training after convergence to $\geq$98success (averaged over the 10 most recent evaluations), or after a maximal training length set per task. The resulting average training length across all 25 tasks was 600k environment steps for VLC MW50 and 750k for the Open X trained model. The corresponding GPU hours vary slightly based on exact architecture used, but we obtained approximately 750k steps in 24 hours on a single NVIDIA V100 or P100 GPU. Our total computational budget was therefore 19.2 GPU hours $\times$ 5 random seeds $\times$ 25 tasks = 2,400 GPU hours for the RL training experiments with VLC MW50 and 24 $\times$ 5 $\times$ 25 = 3,000 hours with VLC Open X. For the MW40 task generalization experiments, the average training length was again 600k for VLC, but 10 random seeds were used, so the corresponding cost for training on 6 tasks was approximately 1150 GPU hours for our method, and slightly more for baselines that took longer to train.
Appendix C Meta-World environment definitions
Appendix D Baseline implementations
For the LOReL baseline experiment, we used the proposed binary classification objective, reversed negatives and 2-frame conditioning (first and last), while keeping all of our other training details and data augmentations identical to our method. This is to ensure LOReL’s smaller and older original architecture did not account for any difference in performance. The original implementations were used for all other methods, and pretrained checkpoints (e.g., HowTo100M [47] pretraining for RoboCLIP) were reused when applicable.
@@ Appendix D Baseline implementations
Appendix E Robustness to visual variations
To evaluate VLC’s generalization to visual observation shift and noise, we run additional experiments where the brightness of the scene is reduced, and uniformly sampled pixel noise is added per image. Example images of this perturbation compared to the original appearance are included in Fig. 4. The success rate curves are included in Fig. 5. 2/6 tasks’ performance is somewhat affected by this noise, but 4/6 are unaffected.
Figure 4: Left: Example observation from the Open Faucet task with the original brightness (0.3) and without image noise. Right: Example observation from the same scene with shifted brightness (0.01) and pixel-wise noise sampled uniformly from [0, 0.2].
Figure 5: Success rates in VLC training under brightness shift and image noise (as shown in Fig. 4).
Appendix F Qualitative evaluation
Appendix G Model-based evaluation results per task
Table 4: Model-based experiment: Accuracy (%) of VLC-MW40 in identifying a successful execution out of 6 sampled trajectories, averaged over 50 random scene initializations for each held-out task. Two action primitives are required to solve Assembly and Pick Out of Hole: first grasping the required object, then moving it as specified, such as lifting it out of the hole.
VLC 100 20 96 100 40 94 54 82 100 100 97 100 80
RoboCLIP 88 60 58 4 86 32 4 14 26 82 44 50 46
LOReL 26 14 44 14 4 22 22 28 72 54 100 30 36